# A Decade of Privacy-Relevant Android App Reviews: Large Scale Trends

### 04 Mar 2024 v1

@@ 6.4 Nigeria
Issues associated with Spying Concerns include Spying on Spouse, Spying on Partner, Allows Unauthorized Calls Access, and Allows Unauthorized Messages Access, suggesting these ads may primarily be targeted to intimate partner abuse perpetrators.
Notably, many of these review ads use similarly formatted email addresses. Although we identified this pattern in privacy reviews, a simple regular expression [^5] Omitted here due to potential harm; contact the authors for information. across all reviews yielded more than 10K matches. A random sample of 100 matches yielded only three false positives. Matching reviews appear on apps with developer-specified categories such as Lifestyle, Tools, Books, Communication, and Dating. Most are from Nigeria. We disclosed this issue to Google Play, and these reviews were taken down.
A simple web search for some of the email addresses from these reviews reveals similar ads posted as comments/reviews across the web as well as posts about the (in)effectiveness of the advertised spying [87], suggesting these services are not unique to Google Play, and may be of interest to the digital intimate partner violence prevention and support community [22, 32, 5].
Finding 7:xxWe detected a subset of reviews offering spyware services, primarily from Nigeria. These were reported to Google Play and have been removed.
@@ 7.1 App types with negative emotions
Social media
These apps receive the second-most privacy reviews (775K) of all app types, and $\sim$57of these reviews are associated with negative emotions. Reviewers most frequently bring up Invasion of Privacy and Data Deletion. Non-specific comments, such as in Invasion of Privacy, include for example: “You invade the privacy of people. Our privacy is sold for profits. You let fake news be posted by foreign countries.” In Data Deletion reviews, users complain that they are unable to delete accounts, posted content, or the social media app itself (some social media apps come pre-installed on phones and are not removable [24]). This latter feedback is more actionable for developers; for example, improving the UI to make data deletion controls more discoverable could potentially help. The strongly negative emotions indicate users who are very frustrated with privacy properties of social apps.
Pet simulators
@@
Child monitors and location tracking
@@ 7.2 Privacy Protective Apps: Quite Positive!
Below we summarize reviews from nine privacy- and security-focused app types. Their dominant privacy themes are shown in Figure 11, and Figure 10 shows the emotion distribution. At a high level, users appreciate the (perceived) protections they receive from these apps. Our results confirm and extend prior work while also highlighting research gaps.
Figure 11: Top security & privacy app types themes (PTPR).
@@
Researchers have found some app lockers may be easily circumventable [51], suggesting that users may be less well protected than they believe. Other research has touched on app lockers’ resemblance to malware [6, 74]. However, despite their wide use, to our knowledge there is little research formally analyzing the security, privacy, and usability of these apps. Although privacy reviews for these apps are decreasing in frequency (§ 5.1), we argue these apps remain an interesting target for future research.
@@
The Photo tools app type has 208K privacy reviews, while Album management has 145K. Reviews for these app types express ‘Joy’ in $>$45% of cases, often praising Photo Protection (e.g., “I love this app. Photos/videos are ALWAYS recorded and protected”) and Content Hiding (e.g., “Greatest application for hiding photos or vids”) features. Unlike app lockers and hiders, these apps allow users to secure media beyond the initial lock screen.
Diary apps
@@
Antivirus The app type has 106K privacy reviews, of which 41% express ‘Joy.’ Password Protection, Privacy Concerns, App Locking, Personal Information Privacy, and Data Deletion are the most common themes. Users comment about a variety of issues outside the traditional antivirus app functionality, which we posit relates to antivirus apps expanding their feature offerings [77]. For example, we see praise for app locking features not only in apps categorized as app lockers, but also in general antivirus apps, many of which offer such features. We see relatively fewer mentions of malware (the primary purpose of antivirus software); examples from issues like Phone Protection and Malware Protection, include: “This is a perfect app. It finds and removes malware very fast. Further, it detects apps that are a threat to your privacy. I 100% recommend it.
The Task & app management app type has 51k privacy reviews, of which 35% are privacy positive. These apps exhibit similar trends to antivirus apps, perhaps because of large overlap in functionality: both managers and antivirus programs offer security and performance features. However, reviews of task management apps include more reviews about Excessive permissions. Though both app types require sensitive permissions to function (e.g., accessing all files), task & app management users may be more hesitant to grant these permissions outside a security p context: “what is up with these crazy permissions? thanks. nope.
@@ 8 Discussion & Conclusion
Our analysis can also be leveraged to identify anomalous behaviors, such as countries with unique patterns of topic discussions (e.g., Nigeria and Türkiye), or countries with abnormally high privacy-review growth rates (e.g., Indonesia). We uncovered large groups of reviewers who use reviews to communicate with developers in ways unlikely to achieve their goals [41]. Examples include ineffective quasi-legal privacy disclaimers in Türkiye, as well as data deletion requests for specific accounts in Indonesia.
@@
We argue that our analysis approach can be useful for a variety of research purposes. We have shown in this paper examples of: (1) corroborating existing research, such as the broad public support for VPNs [17] (in § 7.2), and the divergent views about child monitoring apps ( [15, 88]); (2) adding context to prior work, such as showing that privacy concerns about financial loan apps from Kenya  [57] are being similarly reported in Indonesia, India, and Thailand (in  § 5.3); and (3) identifying unexpected or emerging privacy issues that can seed future work using interviews (to enable direct discussion with users) or surveys (that can direct users to focus on specific aspects). Examples include examining if diary apps and app lockers actually deliver the privacy features offered, and whether users correctly understand the privacy offerings of those apps, in addition to exploring why permissions are a decreasing topic of discussion.
@@ Appendix B Privacy Classifier Baselines
We investigated various modern transformer architectures (excluding traditional models, such as BiLSTM, which are inferior), and selected six models based on comparisons made within the recently published and extensively cited DeBERTaV3 paper [33]. In addition, we also considered BART, T5, and FLAN-T5 models for their popularity. We chose the largest available variant of DeBERTaV3, and chose other models’ sizes accordingly. We use Hark’s [30] T5-11B model as a comparative baseline, our T5-11B model only differs in the diverse training dataset used (§ 3.1). This list of models is not exhaustive, but we believe it offers a reasonable baseline.
@@
In addition, we also experimented with Gemini 1.0 Pro, a state-of-the-art (closed-source) LLM [42]. Despite our prompt engineering (i.e., multiple prompt variants, using few-shot examples.), we could not better the performance T5-11B. Our prompts described our privacy taxonomy, where each high level concept was defined using fine-grained aspects. The prompts further included instances of reviews and their labels as few-shot examples. For the best performing prompt (see Appendix E), the Gemini Pro performed well on recall (0.83) but precision suffered (0.77), with a ROC-AUC of 0.87. We hypothesize that the poor precision arises because the privacy nuances captured in our carefully annotated training dataset are hard to establish using constrained prompts. Exploring LoRA finetuning techniques [36] to finetune an LLM for this task is outside the scope of this paper.
@@
Appendix C PTPR for each Theme per Country.
Figure 12: (Top) Dendrogram of hierarchical clustering of countries based on vectors visualized below. (Bottom) Percentage of a theme within the privacy reviews from each country. Darkest color indicates 10% or more. Lightest color indicates 0%. Both graphs share the same order of countries (x axis).
@@ Appendix D Expanded Privacy Taxonomy
Table 5: Expanded Privacy Taxonomy
Appendix E Privacy Classifier Prompt

### 15 Mar 2024 v3

@@ 6.4 Nigeria
Issues associated with Spying Concerns include Spying on Spouse, Spying on Partner, Allows Unauthorized Calls Access, and Allows Unauthorized Messages Access, suggesting these ads may primarily be targeted to intimate partner abuse perpetrators. Notably, many of these review ads use similarly formatted email addresses. Although we identified this pattern in privacy reviews, a simple regular expression [^5] Omitted here due to potential harm; contact the authors for information. across all reviews yielded more than 10K matches. A random sample of 100 matches yielded only three false positives. Matching reviews appear on apps with developer-specified categories such as Lifestyle, Tools, Books, Communication, and Dating. Most are from Nigeria. We disclosed this issue to Google Play, and these reviews were taken down.
A simple web search for some of the email addresses from these reviews reveals similar ads posted as comments/reviews across the web as well as posts about the (in)effectiveness of the advertised spying[94], suggesting these services are not unique to Google Play, and may be of interest to the digital intimate partner violence research community [22, 34, 5].
Finding 7:xxWe detected a subset of reviews offering spyware services, primarily from Nigeria. These were reported to Google Play and have been removed before publication.
@@ 7.1 App types with negative emotions
Social media:
These apps receive the second-most privacy reviews (775K) of all app types, and $\sim$57of these reviews are associated with negative emotions. Reviewers most frequently bring up Invasion of Privacy and Data Deletion. Non-specific comments, such as in Invasion of Privacy, include: “You invade the privacy of people. Our privacy is sold for profits. You let fake news be posted by foreign countries.” In Data Deletion reviews, users complain that they are unable to delete accounts, posted content, or the app itself (some social media apps come pre-installed and are not removable [24]). This latter feedback is more actionable for developers; for example, improving the UI to make data deletion controls more discoverable could help. The strongly negative emotions indicate users who are very frustrated with privacy properties of social apps.
Pet simulators:
@@
Figure 10: Top security & privacy app types themes (PTPR).
Child monitors and location tracking:
@@ 7.2 Privacy Protective Apps: Quite Positive!
Below we summarize reviews from these privacy- and security-focused app types. Their dominant privacy themes are shown in Figure 10, and Figure 11 shows the emotion distribution. At a high level, users appreciate the (perceived) protections they receive from these apps. Our results confirm and extend prior work while also highlighting research gaps.
Figure 11: Top security & privacy app types with highly positive emotions.
@@
Password Managers:
Out of 27K privacy reviews, reviewers most frequently praise password managers’ main function, Password protection, as well as using biometrics to unlock password storage (Fingerprint Matters). Overall, Password Managers are more frequently labeled with ‘Joy’ (67.2%, compared to 5.6% of ‘Anger’). Accordingly, we see nearly no complaints in top themes. A minority complain that fingerprint authentication doesn’t work, and a smaller minority say they are unable to change their master password.
These complaints don’t necessarily align with problems identified in prior work [71, 78, 57], and we observe few complaints about usability or interaction problems [39]. This discrepancy may be because reviewers likely use password managers, while prior work has often focused on adoption and therefore included participants who do not (yet) use them.
Authenticators:
Despite receiving fewer reviews (11K), authenticators share many similarities with password managers. Reviewers acknowledged these apps’ roles in Password Protection: keeping their Accounts and Personal Information Protected. Reviewers particularly appreciate using fingerprints to approve requests: “Perfect application, just using your finger print you can authenticate apps.” In contrast, a minority express concern that their accounts have been compromised despite using the authenticator, or that they cannot access their accounts due to authenticator problems. These problems are similar to, but more frequent than, issues reported for password managers. Like password managers, authenticators create a single point of authentication failure. Likely due to these issues, users expressed much less joy (44.2%) and much more anger (11.7%) compared to password managers (joy: 67.2%, anger: 5.6%).
User concerns about losing account access have also been identified in prior work [69]. Our analysis demonstrates that these concerns are realized, some of which could arise from lack of adequate instructional material [26]. Some work has focused on security and privacy weaknesses of authenticator implementations [70, 29]. However, these weaknesses, by and large, are invisible to users and do not appear in reviews.
@@
Researchers have found some app lockers may be easily circumventable [53], suggesting that users may be less protected than they believe. Other research has shown app lockers’ resemblance to malware [6, 81]. However, despite wide use, there is little research analyzing the security, privacy, and usability of these apps. We argue these apps remain an interesting research topic.
@@
The Photo tools app type has 208K privacy reviews, while Album management has 145K. Reviews for these app types frequently express ‘Joy’ ($>$45%), often praising Photo Protection (e.g., “I love this app. Photos are ALWAYS protected”) and Content Hiding (e.g., “Greatest application for hiding photos or vids”) features. Unlike app lockers and hiders, these apps allow users to secure media beyond the initial lock screen.
Diary apps:
@@
Antivirus app type has 106K privacy reviews, of which 41% express ‘Joy.’ Password Protection, Privacy Concerns, App Locking, Personal Information Privacy, and Data Deletion are the most common themes. Users comment about a variety of issues outside the traditional antivirus app functionality, which we posit relates to antivirus apps expanding their feature offerings[84]. For example, we see praise for app locking features not only in apps categorized as app lockers, but also in general antivirus apps, many of which offer such features. We see relatively fewer mentions of malware (the primary purpose of antivirus software); examples from issues like Phone Protection and Malware Protection, include: “This is a perfect app. It finds and removes malware very fast. Further, it detects apps that are a threat to your privacy. I 100% recommend it.
The Task & app management app type has 51k privacy reviews, of which 35% are privacy positive. These apps exhibit similar trends to antivirus apps, perhaps because of large overlap in functionality: both managers and antivirus programs offer security and performance features. However, reviews of task management apps include more reviews about Excessive permissions. Though both app types require sensitive permissions to function (e.g., accessing all files), task & app management users may be more hesitant to grant these permissions outside a security context: “what is up with these crazy permissions? thanks. nope.
@@ 8 Discussion & Conclusion
Our analysis can also be leveraged to identify anomalous behaviors, such as countries with unique patterns of topic discussions (e.g., Nigeria and Türkiye), or countries with abnormally high privacy-review growth rates (e.g., Indonesia). We uncovered large groups of reviewers who use reviews to communicate with developers in ways unlikely to achieve their goals [google2023policy]. Examples include ineffective quasi-legal privacy disclaimers in Türkiye, as well as data deletion requests for specific accounts in Indonesia.
@@
We argue that our analysis approach can be useful for a variety of research purposes. We have shown in this paper examples of: (1) corroborating existing research, such as the broad public support for VPNs [17] (in § 7.2), and the divergent views about child monitoring apps [15, 95]; (2) adding context to prior work, such as showing that privacy concerns about financial loan apps from Kenya  [60] are being similarly reported in Indonesia, India, and Thailand (in  § 5.3); and (3) identifying unexpected or emerging privacy issues that can seed future work using interviews (to enable direct discussion with users) or surveys (that can direct users to focus on specific aspects). Examples include examining if diary apps and app lockers actually deliver the privacy features offered, and whether users correctly understand the privacy offerings of those apps, in addition to exploring why permissions are a decreasing topic of discussion.
@@
9 Acknoledgements
We would like to thank Kurt Thomas and Patrick Gage Kelley for their insightful comments on earlier manuscripts. We also thank the anonymous USENIX reviewers for their feedback, and the Google Play team for providing data access and removing spying reviews we identified.
@@ Appendix B Privacy Classifier Baselines
We investigated various modern transformer architectures (excluding traditional models, such as BiLSTM, which are inferior), and selected six models based on comparisons made within the recently published and extensively cited DeBERTaV3 paper [35]. In addition, we also considered BART, T5, and FLAN-T5 models for their popularity. We chose the largest available variant of DeBERTaV3, and chose other models’ sizes accordingly. We use Hark’s [32] T5-11B model as a comparative baseline; our T5-11B model only differs in the diverse training dataset used (§ 3.1). This list of models is not exhaustive, but we believe it offers a reasonable baseline.
@@
In addition, we also experimented with Gemini 1.0 Pro, a state-of-the-art (closed-source) LLM [44]. Despite our prompt engineering (i.e., trying multiple prompt variants and using few-shot examples), we could not better the performance T5-11B. Our prompts described our privacy taxonomy, where each high level concept was defined using fine-grained aspects. The prompts further included instances of reviews and their labels as few-shot examples. For the best performing prompt (see Appendix E), the Gemini Pro performed well on recall (0.83) but precision suffered (0.77), with a ROC-AUC of 0.87. We hypothesize that the poor precision arises because the privacy nuances captured in our carefully annotated training dataset are hard to establish using constrained prompts. Exploring LoRA techniques [38] to finetune an LLM for this task is outside the scope of this paper.
@@
Appendix C Heatmap Showing Theme Distribution per Country.
Figure 12: (Top) Dendrogram of hierarchical clustering of countries based vectors visualized below. (Bottom) Percentage of a theme within the privacy reviews from each country. Darkest color indicates 10% or more. Lightest color indicates 0%. Both graphs share the same order of countries (x axis).
@@ Appendix D Expanded Privacy Taxonomy
Table 5: Privacy taxonomy. Takes [32] (derived from academic taxonomies) and expands per material found in reviews (expert manual analysis). Note this does not represent any formal corporate taxonomy.
Appendix E Privacy Classifier Prompt (derived from Appendix D)